# Qffusion: Controllable Portrait Video Editing via Quadrant-Grid Attention Learning

### v1

@@ Abstract
This paper presents Qffusion, a dual-frame-guided framework for portrait video editing. Specifically, we consider a design principle of “animation for editing”, and train Qffusion as a general animation framework from two still reference images while we can use it for portrait video editing easily by applying modified start and end frames as references during inference. Leveraging the powerful generative power of Stable Diffusion, we propose a Quadrant-grid Arrangement (QGA) scheme for latent re-arrangement, which arranges the latent codes of two reference images and that of four facial conditions into a four-grid fashion, separately. Then, we fuse features of these two modalities and use self-attention for both appearance and temporal learning, where representations at different times are jointly modeled under QGA. Our Qffusion can achieve stable video editing without additional networks or complex training stages, where only the input format of Stable Diffusion is modified. Further, we propose a Quadrant-grid Propagation (QGP) inference strategy, which enjoys a unique advantage on stable arbitrary-length video generation by processing reference and condition frames recursively. Through extensive experiments, Qffusion consistently outperforms state-of-the-art techniques on portrait video editing.
Figure 1: We present Qffusion, a simple yet effective dual-frame-guided portrait video editing framework. Specifically, our Qffusion is trained as a general animation framework from two still reference images whereas it can perform portrait video editing effortlessly when using modified start and end video frames as references during inference. That is, we specify editing requirements by modifying two video frames rather than text. Our Qffusion can perform fine-grained local editing (e.g., modifying age, makeup, hair, style, and sunglasses).
@@ 1 Introduction
Generally, the techniques for portrait video editing can be classified into two categories. One category is based on Generative Adversarial Networks (GANs) [17, 40]. While these methods provide decent processing speeds, they often fail to generalize to unseen humans and suffer from unstable training [54]. The other category comprises diffusion-based methods [20, 47]. These models [53, 10, 41, 16, 28, 27] are pretrained on the large image and video datasets, making them easily model arbitrary objects. Despite the generalizability and controllability of these diffusion models, they still face several challenges, which are listed in Tab. 1. (i) The prevailing text-driven video editing methods[53, 10, 41, 16, 28] struggle to deal with specific local manipulations (e.g., hair editing) since text-driven editing cannot capture sufficient editing details directly. (ii) The very recent method AnyV2V [27] performs first-frame-guided editing, which first employs an image editing model [7, 59] to modify the first frame and then utilizes an image-to-video (I2V) generation model  [62] to propagate such modifications. However, it often leads to a degraded quality, as a single edited frame cannot enforce sufficient appearance coherence. (iii) The video length of the existing methods is always constrained by limited computational resources.
To handle the above challenges, this paper proposes a dual-frame-guided portrait video editing method dubbed Qffusion, which allows for fine-grained local editing on arbitrary-long videos. Specifically, we consider an “animation for editing” design principle and train Qffusion as a video animation network from two still reference images, which can perform portrait video editing effortlessly when applying edited start and end video frames as references during inference. That is, our Qffusion specifies editing requirements by modifying two video frames rather than text or one single frame. For example, we can use professional software (e.g., Photoshop or Meitu11https://www.meitu.com/) for controllable and consistent reference frame editing.
@@
As shown in Fig. 1, Qffusion delivers impressive results in fine-grained local editing, e.g., adding sunglasses, editing age, hair, and style. Besides, since Qffusion is trained as a general video animation framework, we can flexiblely use it for other applications, e.g., whole-body driving [22] and jump cut smooth [51]. Our main contributions are:
  • We propose a Quadrant-grid Arrangement (QGA) scheme to re-arrange reference images and driving signals under a four-grid fashion respectively, which models appearance correspondence and temporal clues all at once.
@@ 2 Related Works
Diffusion Model for Image Generation and Editing. Recently, diffusion models [20, 47] have emerged as a popular paradigm for text-to-image (T2I). DALLE-2 [43] and Imagen [45] can generate high-resolution images via cascaded diffusion models. Then, Stable Diffusion [44] proposes to train diffusion models in the learned latent space for less computational complexity. As for image editing, the early techniques [15, 38, 43, 3, 36] need an editing mask provided by the user, which is time-consuming. To deal with this, there is a line of research conducts text-only image editing [19, 49, 13, 37, 9, 4, 35], which changes the visual content of the input image following the target prompt without masks. Moreover, InstructPix2Pix [7] and MagicBrush [59] perform editing following human instructions. To further improve spatial controllability, ControlNet [60] introduces a side path to Stable Diffusion to accept extra conditions like edges, depth, and human pose.
Diffusion Model for Video Generation and Editing. There have also been substantial efforts in Text-to-Video (T2V). Besides operating diffusion process directly on pixel space [21, 46], the recent T2V models [6, 64, 14, 50] draw inspiration from Stable Diffusion [44] and generate high-quality videos via a learned latent space. Apart from T2V, several representative works [5, 1, 8, 2, 62] lay the cornerstone for image-to-video (I2V) generation.
Regarding text-driven video editing, Tune-A-Video [53] first proposes an efficient one-shot tuning strategy based on Stable Diffusion. Then, a group of methods [16, 28, 10, 41, 11, 56] conduct zero-shot video editing, where various attention mechanisms are designed to capture temporal cues without extra training. Further, following the substantial efforts of I2V generation, AnyV2V [27] conducts image-driven video editing by propagating modified content from the first edited frame. However, it leads to a degraded quality, where those frames that are far away from the first frame usually present an unpleasing reconstruction and editing.
Generally, text-driven video editing methods struggle with certain fine-grained local manipulations on portrait videos, e.g., modifying hair. Different from the typical text-driven video editing methods, Codef [39] introduces a new type of video representation, which consists of a canonical content field and a temporal deformation field recording static contents and transformations separately. By editing the canonical image, Codef can carry out fine-grained local editing. However, it needs training on each video to be edited, whereas ours is a general framework after training.
@@ 3 Preliminary of Stable Diffusion
Stable Diffusion [44] consists of a VAE [26], a diffusion process, and a denoising process. Here, VAE provides a learnable latent space, avoiding the massive resources required for pixel-level calculation.
@@ 4 Methods
We propose a dual-frame-guided portrait video editing method dubbed Qffusion, which can perform fine-grained or local editing on arbitrary-long videos. Specifically, we consider an “animation for editing” principle, and train Qffusion as a video animation framework from two still reference images while we can use it for portrait video editing easily by applying edited start and end video frames as references during inference. That is, we first specify editing requirements by modifying the start and end frames with professional software and then use Qffusion to propagate these fine-grained local modifications to the entire video.
@@ 4.1 Overview
The overall pipeline of our Qffusion is illustrated in Fig. 2. Like SD, our model consists of two parts, i.e., VAE and latent diffusion. In Qffusion, we propose a QGA scheme to arrange four sequential frames into a large four-grid image, where the upper-right and bottom-left frames are masked for generation. This four-grid image is then stacked with their corresponding four-grid keypoints, forming a composite input that encodes both visual and motion information. Formally, given four sequential frames $\{\mathbf{I}^{a},\mathbf{I}^{b},\mathbf{I}^{c},\mathbf{I}^{d}\}$ and their corresponding condition image (i.e., keypoints) $\{\mathbf{C}^{a},\mathbf{C}^{b},\mathbf{C}^{c},\mathbf{C}^{d}\}$, we replace $\mathbf{I}^{b}$ and $\mathbf{I}^{c}$ with all-zero masks and train our model to reconstruct them. The VAE encoder $\mathcal{E}$ first encodes $\mathbf{I}^{a},\mathbf{I}^{d},\mathbf{C}^{a},\mathbf{C}^{b},\mathbf{C}^{c},\mathbf{C}^{d}$ into latent codes $\mathbf{r}^{a},\mathbf{r}^{d},\mathbf{c}^{a},\mathbf{c}^{b},\mathbf{c}^{c},\mathbf{c}^{c}$, respectively. Then, these input latent codes are combined with a noise map to form a fused code through our QGA scheme. Next, a denoiser $\mathbf{\epsilon}_{\theta}$ learns driving correspondence and temporal clues from these latent codes and predicts the denoised latent. Finally, a VAE decoder $\mathcal{D}$ decodes the denoised latent into images. The computation process of Qffusion is:
@@ 4.2 Quadrant-grid Arrangement
Based on SD, we train Qffusion as a general video animation framework from two reference images and four driving signals. Specifically, we propose a Quadrant-grid Arrangement (QGA) scheme to establish the correspondence between two modalities (i.e., appearance features and driving signals) for appearance consistency in the denoiser UNet $\mathbf{\epsilon}_{\theta}$, while jointly modeling the temporal clues at different times.
QGA arranges the latent codes of reference images ($\mathbf{r}^{a}$ and $\mathbf{r}^{d}$) and two all-zero placeholder masks for intermediate frames into a big four-grid image $\{\mathbf{r}^{a},\mathbf{0},\mathbf{0},\mathbf{r}^{d}\}$, where reference images are assigned to upper-left and bottom-right locations. Similarly, we combine four driving conditions into a big four-grid condition image as $\{\mathbf{c}^{a}$, $\mathbf{c}^{b}$, $\mathbf{c}^{c},\mathbf{c}^{d}\}$, which is stacked with the previous four-grid appearance latents $\{\mathbf{r}^{a},\mathbf{0},\mathbf{0},\mathbf{r}^{d}\}$. Here, a one-to-one correspondence between the appearance and conditions is achieved. In this way, appearance representations of different frames would establish spatial relationships in the self-attention layers for the reconstruction of $\tilde{\mathbf{I}}^{b}$ and $\tilde{\mathbf{I}}^{c}$. Besides, temporal clues are also modeled naturally in QGA since motion information is embraced in the composed four-grid representation.
@@
Next, the fused $\mathcal{Q}_{t}$ is fed to denoiser $\mathbf{\epsilon}_{\theta}$ to predict the denoised latent codes. During training, the diffusion process iteratively adds noises to $\mathcal{Z}_{0}$ and eventually leads to $\mathcal{Z}_{T}$. In the denoising process, the denoiser $\mathbf{\epsilon}_{\theta}$ aims to recover latent codes $[[\mathbf{z}^{a},\mathbf{z}^{b}]^{\top},[\mathbf{z}^{c},\mathbf{z}^{d}]^{\top}]$ based on the fused latent $\mathcal{Q}_{T}$.
@@ 4.3 Quadrant-grid Propagation for Inference
The remaining problem is how to continuously generate all intermediate frames given the start and end reference frames and driving signals during inference. Considering our quadrant-grid training design, we also maintain this fashion for inference. Specifically, we only generate two frames for each inference time. Then, we can generate arbitrary-length videos easily by replacing the intermediate conditions sequentially. This makes portrait video editing no longer constrained by limited computing resources.
@@
Figure 4: Animation comparisons with other methods.
@@ 5.1 Implementation Details
5.2 Qualitative Comparison on Animation
Since Qffusion is trained as a video animation framework, we first present a visual comparison with three portrait video animation methods in Fig. 4, which are: (i) GAN. A UNet-based GAN [32], which is trained for reconstruction. Here, two reference images and condition images are used as input to predict corresponding animation images. (ii) ControlNet. We apply ControlNet [60] to encode the condition images and the reference ones. Specifically, we first use two ”reference-only” ControlNet2 2https://github.com/Mikubill/sd-webui-controlnet/discussions/1236 to encode two reference images, and an ”OpenPose” ControlNet3  3https://huggingface.co/lllyasviel/sd-controlnet-openpose for condition encoding. Then, all encoded features are fed into SD to generate animated images that follow the condition motion and reference appearance. (iii) ControlNet $+$ AnimateDiff. We insert the temporal module of AnimateDiff [18] into ControlNet for temporal consistency.
Although Qffusion can generate arbitrary-length videos, we use all methods to generate 80 frames for comparison here. Fig. 4 shows the animation performance, where two reference images are omitted for simplicity. Our Qffusion has the most consistent appearance details and motion. Note that AnimateDiff only generates 16-32 frames. To generate 80 frames, we use an overlap generation strategy (i.e., overlapping 8 frames for a 16-frame generation) to maintain continuity.
Figure 5: Qualitative ablation studies of input arrangement. (a) two images $\{\mathbf{r}^{a},\mathbf{0}\}$; (b) a four-grid arrangement with one reference image $\{\mathbf{r}^{a},\mathbf{0},\mathbf{0},\mathbf{0}\}$ ; (c) our quadrant-grid design $\{\mathbf{r}^{a},\mathbf{0},\mathbf{0},\mathbf{r}^{d}\}$; (d) ground-truth, respectively.

### v2

@@ Abstract
This paper presents Qffusion, a dual-frame-guided framework for portrait video editing. Specifically, we consider a design principle of “animation for editing”, and train Qffusion as a general animation framework from two still reference images while we can use it for portrait video editing easily by applying modified start and end frames as references during inference. Leveraging the powerful generative power of Stable Diffusion, we propose a Quadrant-grid Arrangement (QGA) scheme for latent re-arrangement, which arranges the latent codes of two reference images and that of four facial conditions into a four-grid fashion, separately. Then, we fuse features of these two modalities and use self-attention for both appearance and temporal learning, where representations at different times are jointly modeled under QGA. Our Qffusion can achieve stable video editing without additional networks or complex training stages, where only the input format of Stable Diffusion is modified. Further, we propose a Quadrant-grid Propagation (QGP) inference strategy, which enjoys a unique advantage on stable arbitrary-length video generation by processing reference and condition frames recursively. Through extensive experiments, Qffusion consistently outperforms state-of-the-art techniques on portrait video editing. Project page: https://qffusion.github.io/page/.
Figure 1: We present Qffusion, a simple yet effective dual-frame-guided portrait video editing framework. Specifically, our Qffusion is trained as a general animation framework from two still reference images whereas it can perform portrait video editing effortlessly when using modified start and end video frames as references during inference. That is, we specify editing requirements by modifying two video frames rather than text. In this way, our Qffusion can perform fine-grained local editing (e.g., modifying age, makeup, hair, style, and wearing sunglasses).
@@ 1 Introduction
Generally, the techniques for portrait video editing can be classified into two categories. One category is based on Generative Adversarial Networks (GANs) [20, 45]. While these methods provide decent processing speeds, they often fail to generalize to unseen humans and suffer from unstable training [61]. The other category comprises diffusion-based methods [24, 54]. These models [60, 11, 47, 19, 33, 31] are pretrained on large image and video datasets, making them easily model arbitrary objects. Despite the generalizability and controllability of these diffusion models, they still face several challenges, which are listed in Tab. 1. (i) The prevailing text-driven video editing methods [60, 11, 47, 19, 33] struggle to deal with specific local manipulations (e.g., hair editing) since text-driven editing cannot capture sufficient editing details directly. (ii) The very recent method AnyV2V [31] performs first-frame-guided editing, which first employs an off-the-shelf image editing model [7, 70] to modify the first frame and then utilizes an image-to-video (I2V) generation model  [73] to propagate such modifications. However, it often leads to a degraded quality, as a single edited frame cannot enforce sufficient appearance coherence. (iii) The video length of the existing methods is always constrained by limited computational resources.
To handle the above challenges, this paper proposes a dual-frame-guided portrait video editing method dubbed Qffusion, which allows for fine-grained local editing on arbitrary-long videos. Given that the existing first-frame-guided editing method [31] faces degraded generation quality, we argue that bi-directional context, rather than strictly causal dependencies, is fundamental for maintaining video quality. Specifically, we consider an “animation for editing” design principle and train Qffusion as a video animation network from two still reference images, which can perform portrait video editing effortlessly when applying edited start and end video frames as references during inference. That is, our Qffusion specifies editing requirements by modifying two video frames rather than text or one single frame. For example, we can use professional software (e.g., Photoshop or Meitu11https://www.meitu.com/) for controllable and consistent reference frame editing.
As shown in Fig. 1, Qffusion delivers impressive results in fine-grained local editing, e.g., adding sunglasses, editing age, hair, and style. Besides, since Qffusion is trained as a general video animation framework, we can flexibly use it for other applications, such as whole-body driving [26] and jump cut smooth [58]. Our main contributions are:
  • We propose a Quadrant-grid Arrangement (QGA) scheme to re-arrange reference images and driving signals under a four-grid fashion separately, which models appearance correspondence and temporal clues all at once.
@@ 2 Related Works
Diffusion Model for Image Generation and Editing. Recently, diffusion models [24, 54] have emerged as a popular paradigm for text-to-image (T2I). DALLE-2 [49] and Imagen [52] can generate high-resolution images via cascaded diffusion models. Then, Stable Diffusion [50] proposes to train diffusion models in the learned latent space for less computational complexity. As for image editing, the early techniques [17, 43, 49, 3, 41] need an editing mask provided by the user, which is time-consuming. To deal with this, there is a line of research conducts text-only image editing [23, 56, 14, 42, 10, 4, 40], which changes the visual content of the input image following the target prompt without masks. Moreover, InstructPix2Pix [7] and MagicBrush [70] perform editing following human instructions. In addition, a group of methods [32, 15, 62, 51, 18, 68] conduct personalized text-guided image editing and synthesize novel renditions of several given subjects in different contexts. To further improve spatial controllability, ControlNet [71] introduces a side path to Stable Diffusion to accept extra conditions like edges, depth, and human pose.
Diffusion Model for Video Generation and Editing. Following image generation and editing diffusion models, there have also been substantial efforts in Text-to-Video (T2V). Besides operating diffusion process directly on pixel space [25, 53], the recent T2V models [6, 75, 16, 57] draw inspiration from Stable Diffusion [50] and generate high-quality videos via a learned latent space. Apart from text-driven video generation, several representative works [5, 1, 8, 2, 73] lay the cornerstone for image-to-video (I2V) generation.
Regarding text-driven video editing, Tune-A-Video [60] first proposes an efficient one-shot tuning strategy based on Stable Diffusion. Then, a group of methods [19, 33, 11, 47, 12, 66] conduct zero-shot video editing, where various attention mechanisms are designed to capture temporal cues without extra training. Further, following the substantial efforts of I2V generation, AnyV2V [31] conducts image-driven video editing by propagating modified content from the first edited frame. However, it leads to a degraded quality, where those frames that are far away from the first frame usually present an unpleasing reconstruction and editing. The very recent method Go-with-the-flow [9] also allows for first-frame-guided editing, which replaces random temporal Gaussianity with correlated warped noise derived from optical flow fields. Although promising performance, it still can only edit limited-frame videos, since the relied I2V model is trained on a limited number of frames. In contrast, our Qffusion can perform arbitrary-long portrait video editing.
Generally, text-driven video editing methods struggle with certain fine-grained local manipulations on portrait videos, e.g., modifying hair. Different from the typical text-driven video editing methods, Codef [44] introduces a new type of video representation, which consists of a canonical content field and a temporal deformation field recording static contents and transformations separately. By editing the canonical image, Codef can carry out fine-grained local editing. However, it needs training on each video to be edited, whereas ours is a general framework once it is finished training.
@@ 3 Preliminary of Stable Diffusion
As a powerful image synthesis model, Stable Diffusion [50] consists of a VAE [30], a diffusion process, and a denoising process. Here, VAE provides a learnable latent space, avoiding the massive resources required for pixel-level calculation.
@@ 4 Methods
This paper proposes a dual-frame-guided portrait video editing method dubbed Qffusion, which can perform fine-grained or local editing on arbitrary-long videos. Specifically, we consider an “animation for editing” principle, and train Qffusion as a video animation framework from two still reference images while we can use it for portrait video editing easily by applying edited start and end video frames as references during inference. That is, we first specify editing requirements by modifying the start and end frames with professional software and then use the proposed Qffusion to propagate these fine-grained local modifications to the entire video.
@@ 4.1 Overview
The pipeline of Qffusion is illustrated in Fig. 2. Like SD, our model consists of two parts, i.e., VAE and latent diffusion. In Qffusion, we propose a QGA scheme to arrange four sequential frames into a large four-grid image, where the upper-right and bottom-left frames are masked for generation. This four-grid images are then stacked with their corresponding four-grid keypoints, forming a composite input that encodes both visual and motion information. Formally, given four sequential frames $\{\mathbf{I}^{a},\mathbf{I}^{b},\mathbf{I}^{c},\mathbf{I}^{d}\}$ and their corresponding condition image (i.e., keypoints) $\{\mathbf{C}^{a},\mathbf{C}^{b},\mathbf{C}^{c},\mathbf{C}^{d}\}$, we replace $\mathbf{I}^{b},\mathbf{I}^{c}$ with all-zero masks and train our model to reconstruct them. The VAE encoder $\mathcal{E}$ first encodes $\mathbf{I}^{a},\mathbf{I}^{d},\mathbf{C}^{a},\mathbf{C}^{b},\mathbf{C}^{c},\mathbf{C}^{d}$ into latent codes $\mathbf{r}^{a},\mathbf{r}^{d},\mathbf{c}^{a},\mathbf{c}^{b},\mathbf{c}^{c},\mathbf{c}^{c}$, respectively. Then, these input latent codes are combined with a noise map to form a fused code through our QGA scheme. Next, a denoiser $\mathbf{\epsilon}_{\theta}$ learns driving correspondence and temporal clues from these latent codes and predicts the denoised latent. Finally, a VAE decoder $\mathcal{D}$ decodes the denoised latent into images corresponding to the input conditions. The process of Qffusion is:
@@ 4.2 Quadrant-grid Arrangement
Based on SD, we train our Qffusion as a general video animation framework from two reference images and four driving signals. Specifically, we propose a Quadrant-grid Arrangement (QGA) scheme to establish the correspondence between two modalities (i.e., appearance features and driving signals) for appearance consistency in the denoiser UNet $\mathbf{\epsilon}_{\theta}$, while jointly modeling the temporal clues at different times.
QGA arranges the latent codes of reference images ($\mathbf{r}^{a}$ and $\mathbf{r}^{d}$) and two all-zero placeholder masks for intermediate frames into a big four-grid image $\{\mathbf{r}^{a},\mathbf{0},\mathbf{0},\mathbf{r}^{d}\}$, where reference images are assigned to upper-left and bottom-right locations. Similarly, we combine four driving conditions into a big four-grid condition image as $\{\mathbf{c}^{a}$, $\mathbf{c}^{b}$, $\mathbf{c}^{c},\mathbf{c}^{d}\}$, which is stacked with the previous four-grid appearance latents $\{\mathbf{r}^{a},\mathbf{0},\mathbf{0},\mathbf{r}^{d}\}$. Here, a one-to-one correspondence between the appearance and conditions is achieved. In this way, appearance representations of different frames would establish spatial relationships in the self-attention layers for the reconstruction of $\tilde{\mathbf{I}}^{b}$ and $\tilde{\mathbf{I}}^{c}$. In addition, temporal clues are also modeled naturally in QGA since motion information is embraced naturally in the composed four-grid representation.
@@
Next, the fused $\mathcal{Q}_{t}$ is fed to the denoiser $\mathbf{\epsilon}_{\theta}$ to predict the denoised latent codes. During training, the diffusion process iteratively adds noises to $\mathcal{Z}_{0}$ and eventually leads to $\mathcal{Z}_{T}$. In the denoising process, the denoiser $\mathbf{\epsilon}_{\theta}$ aims to recover latent codes $[[\mathbf{z}^{a},\mathbf{z}^{b}]^{\top},[\mathbf{z}^{c},\mathbf{z}^{d}]^{\top}]$ based on the fused latent $\mathcal{Q}_{T}$.
@@ 4.3 Quadrant-grid Propagation for Inference
The remaining problem is how to continuously generate all intermediate frames given the start and end reference frames and driving signals during inference. Considering our quadrant-grid training design, we also maintain this fashion for inference. Specifically, we only generate two frames for each inference time. Then, we can generate arbitrary-length videos easily by replacing the intermediate conditions sequentially. This makes portrait video animation and editing no longer constrained by limited computing resources.
@@ 5.1 Implementation Details
Figure 4: Animation comparisons with other methods.
Figure 5: Qualitative ablation studies of input arrangement. (a) two images $\{\mathbf{r}^{a},\mathbf{0}\}$; (b) a four-grid arrangement with one reference image $\{\mathbf{r}^{a},\mathbf{0},\mathbf{0},\mathbf{0}\}$ ; (c) our quadrant-grid design $\{\mathbf{r}^{a},\mathbf{0},\mathbf{0},\mathbf{r}^{d}\}$; (d) ground-truth, respectively. All results are generated using the same training and inference strategies.
5.2 Qualitative Comparison on Animation
Since Qffusion is trained as a video animation framework, we first present a visual comparison with several competitive portrait video animation methods in Fig. 4, which are: (i) GAN. A UNet-based GAN [37], which is trained for reconstruction. Here, two reference images and condition images are used as input to predict corresponding animation images. (ii) ControlNet. We apply ControlNet [71] to encode the condition images and the reference ones. Specifically, we first use two ”reference-only” ControlNet2 2https://github.com/Mikubill/sd-webui-controlnet/discussions/1236 to encode two reference images, and an ”OpenPose” ControlNet3  3https://huggingface.co/lllyasviel/sd-controlnet-openpose for condition encoding. Then, all encoded features are fed into SD to generate animated images that follow the condition motion and reference appearance. (iii) ControlNet $+$ AnimateDiff. We insert the temporal module of AnimateDiff [22] into ControlNet for temporal consistency. (iv) Animate Anyone*. We use the re-produced version [27] of Animate Anyone [26] for portrait video animation, where the official code is not publicly available. The re-produced Animate Anyone proposes a face reenactment method using the facial landmarks of driving video to control the pose of the given source image, and keeping the identity of the source image. This face reenactment model is trained on abundant portrait data. (v) MagicAnimate. We employ MagicAnimate [63] for portrait video animation.
Although Qffusion can generate arbitrary-length videos, we use all methods to generate 80 frames for comparison here. We additionally provide a 502-frame example in our Appendix for a long-term stability test. Fig. 4 shows the animation performance, where two reference images are omitted for simplicity. Our Qffusion has the most consistent appearance details and motion. Note that AnimateDiff only generates 16-32 frames. To generate 80 frames, we use an overlap generation strategy (i.e., overlapping 8 frames for a 16-frame generation) to maintain continuity. In addition, MagicAnimate cannot perform portrait-video animation well. The reason is that the method is trained on whole-body data, which can not be generalized to cross-domain data such as portrait video.